OpenStudy (anonymous):

Solve P = 2(l + w) for l. What are the missing values in the table? P w l 14 2 ? 22 8 ? PLEASE HELPPP

hero (hero):

This is easy

OpenStudy (anonymous):

P = 2l + 2w P - 2w = 2l P-2w ----- = l 2 P - 2w so l = -------- 2 now substitute the given values of P and w and calculate the associated value of l